# Fundamental limits of weak learnability in high-dimensional multi-index models

### v1

@@ 2 Settings and definitions
where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the effective process Equation 7. Furthermore, ${{\bm{M}}}^{t}$ is constrained to lie in $\mathcal{S}^{+}_{p}$ due to Nishimori-identity (section G.3 in (Aubin et al., 2019)). The limiting overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ converge in probability to ${{\bm{M}}}^{t}$.
@@ 3 The trivial subspace ($\alpha_{c}=0$)
The state evolution Equation 8 maps the problem of characterising the computational bottlenecks of first-order methods for high-dimensional Gaussian multi-index models to the study of the deterministic, $p$-dimensional dynamical system ${{\bm{M}}}^{t+1}=F({{\bm{M}}}^{t})$. A starting point is identifying its fixed points and their basins of attraction. In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(0,{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior. With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta_{d}(d^{-\nicefrac{{1}}{{2}}})$ numbers. The asymptotic overlap for an uninformed initial condition is thus identically zero ${{\bm{M}}}^{0}=0$, a null-rank matrix. If ${{\bm{M}}}^{0}=0$ is not a fixed point, then ${{\bm{M}}}^{1}\succ 0$, implying the weak recovery of a subspace of dimension $k={\rm rank}({{\bm{M}}}^{1})>0$ with just a single step of AMP.
@@
${\bm{g}}_{\rm out}(y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y=y]=\mathbf{0},$ (9)
@@
Let $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ denote the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that:
${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|Y=y]=0,$ (10)
almost surely over $Y$ Equation 7. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
Note that for single-index models ($p=1$), the condition in Equation 9 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020b). Interestingly, this is exactly the same condition appearing in Damian et al. (2024b) for weak learnability of single-index models in the SQ model; see eq. (3) therein. To make Lemma 3.1 concrete, let us look at a few examples. A detailed derivation of these examples is discussed in the Appendix C.
@@
$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[(\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y])^{2}\right]\neq 0,$ (12)
where ${\bm{v}}\in T^{\star}$. Indeed, this can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$. The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019). In fact the denoiser ${\bm{g}}_{\rm out}$ is the optimal such transformation in the sense that when ${\bm{g}}_{\rm out}$ fails to obtain a linear correlation along $\mathbf{v}$, i.e when $\mathbf{v}\in H^{\star}$, then no transformation can:
@@
For any $\mathbf{u}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{u}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (13)
Note that this is a non-asymptotic statement about the denoising function ${\bm{g}}_{\rm out}$, and the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
@@
$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}W^{\star}\mathbf{v}=\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$. ∎
This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $O(d)$ samples. The situation can be very different when reusing batches: Dandi et al. (2024b) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above. However, this transformation may not be optimal. For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (11) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024b).
@@ 4 Computational phase transitions ($\alpha_{c}>0$)
with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$. Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is stable a fixed point for any $n=\Theta(d)$.
The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section B.2 for a proof. A consequence of this result is that for $\alpha<\alpha_{c}$, even with infinitesimally small informed initialisation AMP will converge back to the trivial fixed point and fail to recover any subspace, while for $\alpha>\alpha_{c}$, it develops a growing overlap along a non-empty subspace starting with arbitrarily small initialization. This marks the onset of a computational phase transition for weak learnability in the first-order methods class. Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020a)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016). The easy subspace spanned by the directions that can be learned with $n>\alpha_{c}d$ samples can be characterised as follows:
@@
Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank. Then, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1. For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=0$ is asympotically stable i.e. there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}}\rightarrow 0$ as $t\rightarrow\infty$. Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
Heuristically, random initialization is equivalent to setting $\epsilon=O(\nicefrac{{1}}{{\sqrt{d}}})$ in the above theorem. Therefore, we expect that weak-recovery along $E^{\star}$ can be achieved in $O(\log d)$ iterations. As for Lemma 3.1, the expression for the weak recovery threshold Equation 16 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020a). Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore
@@
which is exactly eq. (11) in Barbier et al. (2019). In other words, as long as the link function $g$ has a non-zero second-order coefficient in the Hermite basis, it is learnable by Algorithm 1 with $n>\alpha_{c}d$ samples. Optimality of AMP then implies that no first-order method can achieve non-vanishing correlation with ${\bm{w}}^{\star}$ when $n<\alpha_{c}d$ as $d\to\infty$.
@@
The sparsity parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003). [^3] Strictly speaking, the noiseless parity is efficiently learnable by Gaussian elimination (Blum et al., 1994). Here, we consider computational efficiency in a noise-tolerant sense. Therefore, it should come as no surprise that $\alpha_{c}(p)=\infty$ for $p\geq 3$, providing a concrete example of what a hard subspace (i.e. that cannot be learned with $n=\Theta(d)$ samples) looks like for AMP. Similar to the discussion of trivial directions in Lemma 3.2, we can invoke optimality of AMP to translate our result on the optimal denoiser $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$ to a statement for queries on general label pre-processing transformations:
@@
For any $\mathbf{v}^{\star}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\star},\mathbf{x}\rangle)\right]=0$ (20)
Note that this is a non-asymptotic statement about the denoising function ${\bm{g}}_{\rm out}$, and the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 7).
For an AMP-hard function (that is, when $\alpha_{c}=\infty$), this implies that, even after applying any transformation to the output (as allowed in SQ) an algorithm like SGD would require at least $O(d^{2})$ data, or $O(d^{3/2})$ if we allow smoothing as in Damian et al. (2024a).

### v2

@@ 2 Settings and definitions
where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the following effective process
$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{\xi}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (8)
with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$. Furthermore, ${{\bm{M}}}^{t}$ is constrained to lie in $\mathcal{S}^{+}_{p}$ due to Nishimori-identity (section G.3 in (Aubin et al., 2019)).
As a consequence, the asymptotic mean-squared error on the label prediction is given by:
${\mathbb{E}_{\bm{X},\bm{y}}}\mathbb{E}_{\bm{x},y}\left[\left(y-g\left(\hat{\bm{W}}^{t}(\bm{X},\bm{y})\bm{x}\right)\right)^{2}\right]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (9)
where the expectation is taken over the effective estimation process Equation 8.
@@ 3 The trivial subspace ($\alpha_{c}=0$)
The state evolution Equation 7 maps the problem of characterising the computational bottlenecks of first-order methods for high-dimensional Gaussian multi-index models to the study of the deterministic, $p$-dimensional dynamical system ${{\bm{M}}}^{t+1}=F({{\bm{M}}}^{t})$. A starting point is identifying its fixed points and their basins of attraction. In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(0,{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior. With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta(d^{-\nicefrac{{1}}{{2}}})$ element-wise. The asymptotic overlap for an uninformed initial condition is thus identically zero ${{\bm{M}}}^{0}=0$, a null-rank matrix. If ${{\bm{M}}}^{0}=0$ is not a fixed point, then ${{\bm{M}}}^{1}\succ 0$, implying the weak recovery of a subspace of dimension $k={\rm rank}({{\bm{M}}}^{1})>0$ with just a single step of AMP.
@@
${\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y]=\mathbf{0},$ (10)
@@
We define $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ as the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that the following holds almost surely over $Y=g({\bm{Z}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
${\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|y=Y]=0,$ (11)
where the expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1. Equation 8. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
Note that for single-index models ($p=1$), the condition in Equation 10 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020). Interestingly, this is exactly the same condition appearing in Damian et al. (2024) for weak learnability of single-index models in the SQ model; see eq. (3) therein. To make Lemma 3.1 concrete, let us look at a few examples. A detailed derivation of these examples is discussed in the Appendix C.
@@
$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y]^{2}\right]\neq 0,$ (13)
where ${\bm{v}}\in T^{\star}$. Indeed, the left-hand side can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$. The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019). In fact the denoiser ${\bm{g}}_{\rm out}$ is the optimal such transformation in the sense that when ${\bm{g}}_{\rm out}$ fails to obtain a linear correlation along $\mathbf{v}$, i.e when $\mathbf{v}\in H^{\star}$, then no transformation can:
@@
For any $\mathbf{v}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (14)
Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
@@
$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$. ∎
This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $\mathcal{O}(d)$ samples, as the absence of a first-order Hermite coefficient implies a sample complexity of $\mathcal{O}(d\log d)$ as shown in Theorems 1.3 and 1.4 in Arous et al. (2021). The situation can be very different when reusing batches: Dandi et al. (2024) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above. However, this transformation may not be optimal. For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (12) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024).
@@ 4 Computational phase transitions ($\alpha_{c}>0$)
with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$. Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is a stable fixed point for any $n=\Theta(d)$.
The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section B.2 for a proof. A consequence of this result is that for $\alpha<\alpha_{c}$, even with infinitesimally small informed initialisation AMP will converge back to the trivial fixed point and fail to recover any subspace, while for $\alpha>\alpha_{c}$, it develops a growing overlap along a non-empty subspace starting with arbitrarily small initialization. This marks the onset of a computational phase transition for weak learnability in the first-order methods class. Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016). The easy subspace spanned by the directions that can be learned with $n>\alpha_{c}d$ samples can be characterised as follows:
@@
Let ${{\bm{M}}}_{d}^{t}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ denote the model-target overlap matrix at any finite time $t$. Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank. Then,under Assumption 1, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1. For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=0$ is asympotically stable i.e. there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{d}}\rightarrow 0$ as $t\rightarrow\infty$. Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
Note that in the above the initial overlap scale $\epsilon>0$ can be arbitrarily small but constant with respect to $d$. Heuristically, $\epsilon$ is often identified with $\mathcal{O}(\nicefrac{{1}}{{\sqrt{d}}})$, the scale of a random initial condition. Although this heuristic is not covered by Theorem 4.2, we have numerically checked it is consistent with the extrapolation of the theory, which suggests that for $\alpha>\alpha_{c}$ easy subspaces can be weakly recovered with $\mathcal{O}(\log d)$ steps, see Figure 2.
As for Lemma 3.1, the expression for the weak recovery threshold Equation 17 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020). Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore
@@
which is exactly eq. (11) in Barbier et al. (2019). Optimality of AMP then implies that no first-order method can achieve non-vanishing correlation with ${\bm{w}}^{\star}$ when $n<\alpha_{c}d$ as $d\to\infty$.
@@
The sparse-parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003). [^4] Strictly speaking, the noiseless parity is efficiently learnable by Gaussian elimination (Blum et al., 1994). Here, we consider computational efficiency in a noise-tolerant sense. Therefore, it should come as no surprise that $\alpha_{c}(p)=\infty$ for $p\geq 3$, providing a concrete example of what a hard subspace (i.e. that cannot be learned with $n=\Theta(d)$ samples) looks like for AMP. Similar to the discussion of trivial directions in Lemma 3.2, we can invoke optimality of AMP to translate our result on the optimal denoiser $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$ to a statement for queries on general label pre-processing transformations:
@@
For any $\mathbf{v}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0$ (21)
Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 8).
For an AMP-hard function (that is, when $\alpha_{c}=\infty$), the above Lemma implies that, even after applying any transformation to the output (as allowed in SQ) online-SGD would require at least $\Omega(d^{2})$ data, since online-SGD with squared loss requires $\Omega(d^{\kappa-1})$ samples/iterations for weak-recovery of a target function when the first $\kappa$ Hermite-coefficients are 00 (Theorems 1.3 and 1.4 in Arous et al. (2021)). Furthermore, for single-index models satisfying the condition in Lemma 4.3 (i.e $\alpha_{c}=\infty$), the SQ lower-bound predicts a sample complexity requirement of at-least $\Omega(d^{3/2})$ (Damian et al., 2024). Therefore, learning such functions with $\mathcal{O}(d)$ samples is conjectured to be algorithmically-hard.